OpenStudy (anonymous):

Write a quadratic equation whose solutions are 7i and -7i. The equation is: x^2+____=0

OpenStudy (anonymous):

49

OpenStudy (anonymous):

not much work here just \[7-\times -7i=-49i^2=49\] or think if the zeros are \[7i,-7i\] then you must be solving \[x^2=-49\]

OpenStudy (anonymous):

awesome, thank you